Question 3 a) In the Sunday newspaper, there are usually coupons that you can clip and take to the store to save money on products. Anyone can buy a newspaper, and the value of the coupons easily exceeds the price of the newspaper for most consumers. Is this an example of price discrimination? Explain. (5 marks) b) Differentiate between second degree and third degree price discrimination. Support (7 marks) your answer with suitable examples. c) Georges Pants Store sells a particular brand of slacks. It recently decreased the price of a popular brand by 5 percent. The decrease was followed by a 20 percent increase in sales. The marginal cost of producing these slacks is RM100 i. What is the point price elasticity of demand for these slacks? ii. What would be the optimal markup and price of these items? (5 marks) (8 marks)
Answer #1

Answer 1 a : Yes the above scanerio is an example of price discrimination. Price discrimination means charging different price from different customers. Providing coupon at discounted rate is a result of charging incentives to the customer on Sunday .Here the newspaper discriminate on the basis of the price. They charged different price from the customer as in case of providing incentives.

b : Difference between second degree and third degree price discrimination are :

  1. Second degree price discrimination means charging different price on the amount of Quantity consumed. example : after certain number of unit consumed electricity tariff has been reduced.
  2. Third degree price discrimination means charging different prices from different set of consumer. The form may separate the market on the basis of the consumers.example : income and age group of the consumer.

Answer c : point price elasticity of demand =% change in qty / % change in price = 20%/5%= 4

Optimal mark up price = ( e​​​​​ p / 1+e p) MC

P= ( 4/1+4)*100= (4/5)*100= 80